40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Dec 12 00:47:45 2013, Den, Update, LSC, locking activity POP_ASC.pdf
    Reply  Fri Dec 13 02:19:57 2013, Jenne, Update, LSC, locking activity PRMI_Yarm_NewPRCLasc.pngtry_dc.pdf
       Reply  Fri Dec 13 02:30:22 2013, Koji, Update, LSC, locking activity 
Message ID: 9463     Entry time: Fri Dec 13 02:30:22 2013     In reply to: 9462
Author: Koji 
Type: Update 
Category: LSC 
Subject: locking activity 

According to the measurement by Eric, the X-arm green PDH UGF is too low. We still have some room to increase the gain.

The out of loop stability of the ALS for each arm should be measured everyday.
Otherwise we can't tell whether the arm is prepared for advanced locking activities or not.

We expect to see the arm stablity of ~50pm_rms for the Y arm and ~150pm_rms for the X arm.

ELOG V3.1.3-